OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3.

pooja195 (pooja195):

\[\huge~y-y1=m(x-x1)\]

pooja195 (pooja195):

(x1,y1) m=the slope :) So just plug them in! :) and please list the answer choices

OpenStudy (anonymous):

Ok, one sec

OpenStudy (anonymous):

y-6=-3x -3 y-6=-3(x+1) y=-3x+3 3x+y+3

pooja195 (pooja195):

ok so now just plug in the numbers into the formula i gave u earlier :)

pooja195 (pooja195):

Do you get it?

OpenStudy (anonymous):

Im doing it now :P

pooja195 (pooja195):

:)

OpenStudy (anonymous):

I think its the first one

pooja195 (pooja195):

No quite....remeber we are not distributing they only want it in the point slope form try again :)

OpenStudy (anonymous):

I dont know, I ended up using a graphing calculator and none of them worked?